OpenStudy (anonymous):

Function and Composite OK, is this correct? Find the functions f(x) and g(x) so that f(g(x))=I5x+4I (those are AV lines) f(x)=5x; g(x)=4 Answer 24

OpenStudy (anonymous):

the answer isn't a number, it is two functions, \(f\) and \(g\)

OpenStudy (anonymous):

you have lots of choices, but one could be \[g(x)=5x+4\] and the other \[f(x)=|x|\] that way \[f(g(x))=f(5x+4)=|5x+4|\] as needed

OpenStudy (anonymous):

oh man i am very confused then. how do you know how to break this down?